Dashboard Inc. manufactures and assembles automobile instrument panels for both eCar Motors and Greenville Motors. The process consists of a lean product cell for each customer’s instrument assembly. The data that follow concern only the eCar lean cell.

For the year, Dashboard Inc. budgeted the following costs for the eCar production cell:

Conversion Cost Categories Budget
Labor $800,000
Supplies 275,000
Utilities 325,000
   Total $1,400,000

Dashboard Inc. plans 2,000 hours of production for the eCar cell for the year. The materials cost is $240 per instrument assembly. Each assembly requires 24 minutes of cell assembly time. There was no April 1 inventory for either Raw and In Process Inventory or Finished Goods Inventory.

The following summary events took place in the eCar cell during April:

a. Electronic parts and wiring were purchased to produce 450 instrument assemblies in April.
b. Conversion costs were applied for the production of 400 units in April.
c. 380 units were started, completed, and transferred to finished goods in April.
d. 350 units were shipped to customers at a price of $800 per unit.

Required:

1. Determine the budgeted cell conversion cost per hour.
$ per hour

2. Determine the budgeted cell conversion cost per unit.
$ per unit

3. Journalize the summary transactions (a) through (d). If an amount box does not require an entry, leave it blank.

a.
b.
c.
d. Sale
d. Cost

4. Determine the ending balance in Raw and In Process Inventory and Finished Goods Inventory.

Raw and In Process Inventory: $
Finished Goods Inventory: $
Answer #1
1
Budgeted cell conversion cost per hour 700 per hour =1400000/2000
2
Budgeted cell conversion cost per unit 280 per unit =700*(24/60)
3
Raw and in process inventory 108000 =450*240
     Accounts Payable 108000
Raw and in process inventory 126000 =450*280
       Conversion costs 126000
Finished goods inventory 197600 =380*(240+280)
       Raw and in process inventory 197600
Accounts Receivable 280000 =350*800
       Sales 280000
Cost of goods sold 182000 =350*(240+280)
        Finished goods inventory 182000
4
Raw and In Process Inventory: 36400 =108000+126000-197600
Finished Goods Inventory: 15600 =197600-182000
Answer #2

thank you random nigga above me who answered

answered by: thanks